OpenStudy (anonymous):

Which sentence explains the correct first step in the solution of this equation? 4 • (x – 3) = 9 A) add 2 to both sides B)Apply the distributive property to get 4x - 12=9 C) Apply the distributive property to get 4x - 3=9 D) Subtract 9 from both sides

OpenStudy (internettroll11):

The answer is B. You will distribute 4(x-3)=9 to get 4x-12=9

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

can u help me with another question

OpenStudy (internettroll11):

sure

OpenStudy (anonymous):

Mika’s gym allows her to bring 3 guests for free. After the first three, she must pay $7 per guest. How many guests, g, can she bring to the gym if she pays $42? Three of these equations give the correct value of g. Which equation does NOT? A) 7g = 63 B) 7g - 3 = 42 C) 7g(g- 3) =42 D) g=9

OpenStudy (anonymous):

for C) 7(g - 3)= 42 * i mean

OpenStudy (internettroll11):

g should be 9. so A, D and C are correct values of g. B is not.

OpenStudy (internettroll11):

If shes paying $42 dollars that meant she paid for 6 guests and she can also bring 3 free for a total of 9 ppl.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

can i ask 2 more questions?

OpenStudy (anonymous):

Solve n -4 (n - 6) = 12

OpenStudy (anonymous):

Start by simplifying. -4(n-6)=12 becomes -4n + 24 = 12